# LapSum – One Method to Differentiate Them All: Ranking, Sorting and Top-k Selection

### v1

@@ Abstract
We present a novel technique for constructing differentiable order-type operations, including soft ranking, soft top-k selection, and soft permutations. Our approach leverages an efficient closed-form formula for the inverse of the function $\mathrm{LapSum}$, defined as the sum of Laplace distributions. This formulation ensures low computational and memory complexity in selecting the highest activations, enabling losses and gradients to be computed in $O(n\log{}n)$ time. Through extensive experiments, we demonstrate that our method outperforms state-of-the-art techniques for high-dimensional vectors and large $k$ values. Furthermore, we provide efficient implementations for both CPU and CUDA environments, underscoring the practicality and scalability of our method for large-scale ranking and differentiable ordering problems.
Machine Learning, ICML
@@ 1 Introduction
Neural networks are trained using data through gradient descent, which requires models to be differentiable. However, common ordering tasks such as sorting, ranking, or top-k selection are inherently non-differentiable due to their piecewise constant nature. This typically prevents the direct application of gradient descent, which is essential for efficient learning from data. These challenges have become increasingly relevant in recent years (Lapin et al., 2016; Blondel et al., 2020; Petersen et al., 2022b). To address them, relaxations or approximations are needed to make such tasks compatible with neural network training. This can be achieved by smoothing the objective functions, introducing uncertainty into algorithms, softening constraints, or adding variability. The methods include smoothed approximations (Berrada et al., 2018; Garcin et al., 2022), optimal based (Cuturi et al., 2019; Xie et al., 2020), permutations (Petersen et al., 2022b). Some of these solutions are frequently not closed form.
@@
Challenges in addressing order related tasks arise from several factors. Among them are abrupt changes and instabilities when traversing regions. Time and memory issues often limit the solutions for large $n$ and $k$ values. There are also issues with scalability and batch processing. Techniques employed to solve these problems include relaxations and estimators, ranking regularizers, even learning-based ranking solutions. However, some problems remain unsolved or the existing methods are incomplete. E.g. efficiency in terms of time and memory usage may still be inadequate, impeding the solution of certain tasks (Xie et al., 2020). Frequently, there is a speed-precision trade-off. Several methods lack closed-form solutions, confining one to approximations or complicated calculations. Some methods do not yield probabilities (Blondel et al., 2020). Several still do not support, for example, easy-to-use GPU methods.
Our aim was to address the above-mentioned issues. The main achievement of this paper is the construction of the general theory based on arbitrary density, such that incorporation of the Laplace distribution provides a closed-form, numerically stable, and efficient solutions to all previously mentioned problems. More precisely, our approach, called $\mathrm{LapSum}$, is based on the sum of Laplace distributions, see the scheme Fig. 1 for construction of soft top-k selection problem. We show that this approach gives a solution that is both theoretically sound and practical: the reduced $O(n\log{}n)$ time complexity is on par with that of sequence sorting, while the memory requirements are very limited in comparison to other available top-k solutions.
@@
  • propose a novel, theoretically sound simple closed-form solution, called $\mathrm{LapSum}$ [^1] Our top-k solution’s pseudo-code has 26 lines (Appendix A), with full code will be available on GitHub in the camera-ready. , for all classical soft-order based problems,
  • prove and experimentally verify that $\mathrm{LapSum}$ has $O(n\log{}n)$ low time and $O(n)$ memory complexities, together with derivatives, outperforming within this aspect all other existing methods,
  • offer easy to use code for $\mathrm{LapSum}$, for both CPU and CUDA, making our approach feasible for large optimization problems.
2 Related Works
The strategy of direct top-k training to optimize network performance, rather than the straightforward use of top-1, was first introduced in fast top-k SVM (Lapin et al., 2016). The proposition called for the construction of differentiable loss functions to enable gradient learning.
There are several methodologies available. Some focus on smooth approximations. Berrada et al. (2018) proposed a smooth top-k to more accurately approximate the top-k loss and facilitate a direct technique for top-k gradient learning. While these and comparable methods for direct gradient-based top-k learning yield precise results, they can be computationally intensive (Garcin et al., 2022). Conversely, some smooth approximations can speed up optimization, but may sacrifice accuracy.
@@ 2 Related Works
Multiple studies demonstrate the effective application of top-k learning techniques in pattern recognition and feature extraction, the creation of sparse networks and feature extraction, and recommender systems, such as through the use of differentiable logic gates (Zhao et al., 2019; Hoefler et al., 2021; Chen et al., 2023; Xu et al., 2023; Chen et al., 2021).
@@ 3 General soft-order theory based on sums of cumulative density functions
We begin by demonstrating how a cumulative density function can be used to construct soft analogues of hard, order-based problems such as top-k selection and permutations. While the proposed approach is primarily theoretical and not computationally efficient for a general density, we show in the next section that it can be efficiently computed with complexity $O(n\log n)$ in the specific case of the Laplace distribution.
@@ 3.1 Function ${F\text{-}\mathrm{Sum}}$
We base our construction on an arbitrary even and strictly positive density function $f$, with $F$ denoting its cumulative density function. Note that then $F(-x)=1-F(x)$ and consequently $F(0)=1/2$. Additionally, for a scale parameter $\alpha\neq 0$ we put
@@
${F\text{-}\mathrm{Sum}}_{\alpha}(x;r)=\sum_{i}F_{\alpha}(x-r_{i}).$
Clearly, the image of ${F\text{-}\mathrm{Sum}}_{\alpha}$ is the interval $(0,n)$. It occurs that ${F\text{-}\mathrm{Sum}}_{\alpha}(x;r)$ is invertible as a function of $x$. This follows from the fact that since $f$ is strictly positive density, $F_{\alpha}$ is strictly increasing continuous function for $\alpha>0$ and strictly decreasing for $\alpha<0$. Consequently, ${F\text{-}\mathrm{Sum}}_{\alpha}(x;r)$ is a continuous strictly monotonous function, which can be concluded in the following observation.
@@
Let the sequence $r=(r_{0},\ldots,r_{n-1})\in\mathbb{R}^{n}$ be given, and let $w\in(0,n),\alpha\neq 0$ be arbitrary. Then the equation
${F\text{-}\mathrm{Sum}}_{\alpha}(x;r)=w$
has a unique solution.
@@
It is worth noting that the above functions are differentiable and the derivatives can efficiently be computed. By $f_{\alpha}$ we denote the function $f_{\alpha}(x)=\frac{1}{\alpha}f(x/\alpha)$. Observe that, for $\alpha>0$, it is the density which CDF equal to $F_{\alpha}$. Then we have
$\frac{\partial}{\partial x}{F\text{-}\mathrm{Sum}}_{\alpha}(x;r)=\sum_{i}f_{\alpha}(x-r_{i}).$ (1)
Now we compute the derivatives of the inverse function $w\to{F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(w;r)$. Given $w\in(0,n)$, let $b=b_{\alpha}(w;r)$ denote the unique solution to
${F\text{-}\mathrm{Sum}}(b)=\sum_{i}F_{\alpha}(b-r_{i})=w.$
We want to compute the derivative of the function $b$ with respect to $w$. To do so, we differentiate the above formula with respect to $w$ and obtain: $\frac{\partial b}{\partial w}\sum_{i}f_{\alpha}(b-r_{i})=1$. Consequently
$\frac{\partial}{\partial w}{F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(w;r)=\frac{\partial b}{\partial w}=\frac{1}{\sum_{i}f_{\alpha}(b-r_{i})}.$ (2)
@@ 3.2 Soft rankings and soft orderings
Figure 2: We present the computation of the soft ranking based on $F$ given as CDF of Laplace distribution, here $(s_{0},s_{1},s_{2})={F\text{-}\mathrm{Rank}}_{1}(r_{0},r_{1},r_{2})$.
For $\alpha\neq 0$ we define
${F\text{-}\mathrm{Rank}}_{\alpha}(r_{k})={F\text{-}\mathrm{Sum}}_{\alpha}(r_{k};r)-\tfrac{1}{2},$
see Fig. 2. One can easily observe that ${F\text{-}\mathrm{Rank}}_{\alpha}$ is permutation equivariant i.e., a permutation of the input will result in a similar permutation of the output. By $s^{+}_{i}$ we the rank of $r_{i}$ in the sequence $r$ (with respect to standard order), that is
@@
By $s_{i}^{-}$ we denote the rank of element $r_{i}$ in the reverse ordering of $r$, i.e. $s_{i}^{-}=(n-1)-s_{i}^{+}$. To prove that this is a correct soft version of ranking operation, we need to show that in the limiting case we obtain the hard ranking given by
$\mathrm{rank}^{\pm}(r)=(s_{i}^{\pm})_{i=0..n-1}.$
@@
$\lim_{\alpha\to 0^{\pm}}{F\text{-}\mathrm{Rank}}_{\alpha}(r)=\mathrm{rank}^{\pm}(r).$
@@
Let us prove the limiting formula above for $\alpha>0$ (the case $\alpha<0$ is similar). Since $F_{\alpha}$ (as $\alpha\to 0^{+}$) converges to a Heaviside function with $1/2$ value at zero, we have
$\lim\limits_{\alpha\to 0^{+}}\!{F\text{-}\mathrm{Rank}}_{\alpha}(r_{k})=\!\lim\limits_{\alpha\to 0^{+}}\!\big{(}F_{\alpha}(r_{k})+\!\!\!\sum\limits_{i:r_{i}\neq r_{k}}\!\!F_{\alpha}(r_{k}-r_{i})\big{)}$
$=\tfrac{1}{2}+\!\!\sum\limits_{i:r_{i}\neq r_{k}}\!\!H(r_{k}-r_{i})=\tfrac{1}{2}+s_{k}^{+}.\hskip 70.0001pt\qed$
@@
Soft ordering can be seen intuitively as a partially inverse operator to soft ranking. We define
${F\text{-}\mathrm{Sort}}_{\alpha}(r)={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(\tfrac{1}{2}+(0,1,\ldots,n-1)).$
@@
  • ${F\text{-}\mathrm{Sort}}_{\alpha}(r)$ converges to the sorted $r$ for $\alpha\to 0^{+}$,
  • ${F\text{-}\mathrm{Sort}}_{\alpha}(r)$ converges to $r$ with respect to decresasing order as $\alpha\to 0^{-}$.
Figure 3: We show the result of soft sorting base on $F$ given by CDF of Laplace distribution. Observe that for small $\alpha$ the constructed sequence $(q_{0},q_{1},q_{2})={F\text{-}\mathrm{Sort}}_{1}(r_{0},r_{1},r_{2})$ practically coincides with the sorted sequence $r$.
@@ 3.3 Soft top-k
We proceed to the construction of soft differentiable version of top-k selection problem. The aim is to construct a differentiable version of the hard top-k (formally max top-k) operation, which for a sequence $r=(r_{i})\in\mathbb{R}^{n}$ returns a binary sequence $(p_{i})\in\{0,1\}^{n}$ which has value $1$ for indices with top-k greatest values of $r$, and zero for all other indices. Clearly $\sum_{i}p_{i}=k$.
In our approach we can choose arbitrary $w\in(0,n)$, where we do not require $w$ to be an integer. Using $\mathrm{LapSum}$ we construct the parameterization of the space
$\Delta_{w}=\{[p_{0},\ldots,p_{n-1}]\in(0,1)^{n}:\sum_{i}p_{i}=w\}.$
@@
$\begin{array}[]{c}{F\text{-}\mathrm{Top}}_{\alpha}:(r;w)=(F_{\alpha}(b-r_{0}),\ldots,F_{\alpha}(b-r_{n-1}))\\ \text{ where }b={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w;r),\end{array}$
see Fig. 1 for viusalization of this approach.
@@
For every $w\in(0,n)$ and $r\in\mathbb{R}^{n}$ we have
${F\text{-}\mathrm{Top}}_{\alpha}(r;w)\in\Delta_{w}.$ (3)
Moreover, if $k=w$ is integer and the elements of $r$ are pairwise distinct, then
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r;k)\to\mathop{\mathrm{topmin}}_{k}(r)$ as $\alpha\to 0^{+}$,
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r;k)\to\mathop{\mathrm{topmax}}_{k}(r)$ as $\alpha\to 0^{-}$.
@@
Observe that (3) follows directly from the fact that $b={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w;r)$, and thus $\sum_{i}F_{\alpha}(b-r_{i})=w$.
Due to the limited space the proof of the limiting case is given in Appendix, Theorem B.1. ∎
@@
One can observe that the function ${F\text{-}\mathrm{Top}}$ has properties similar to that of softmax. In particular, if two outputs are equal, then the inputs are equal modulo some constant translation. Namely, by the invertibility of $F$ we obtain that if ${F\text{-}\mathrm{Top}}_{\alpha}(r;w)={F\text{-}\mathrm{Top}}_{\alpha}(\bar{r};w)$ then
@@
where $b={F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(r;w)$, $\bar{b}={F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(\bar{r};w)$. This implies in particular, that similarly to softmax, ${F\text{-}\mathrm{Top}}_{\alpha}$ restricted to $\{r=(r_{i})\in\mathbb{R}^{n}:\sum_{i}r_{i}=0\}$ is invertible.

### v2

@@ Abstract
We present a novel technique for constructing differentiable order-type operations, including soft ranking, soft top-k selection, and soft permutations. Our approach leverages an efficient closed-form formula for the inverse of a function $\mathrm{LapSum}$, defined as a sum of Laplace distributions. This formulation ensures low computational and memory complexity in selecting the highest activations, enabling losses and gradients to be computed in $O(n\log{}n)$ time. Moreover, $\mathrm{LapSum}$ can easily be parallelized, both with respect to time and memory. Through extensive experiments, we demonstrate that our method outperforms state-of-the-art techniques for high-dimensional vectors and large $k$ values. Furthermore, we provide efficient implementations for both CPU and CUDA environments, underscoring the practicality and scalability of our method for large-scale ranking and differentiable ordering problems.
Machine Learning, ICML, top-k learning
@@ 1 Introduction
Neural networks are trained using data through gradient descent, which requires models to be differentiable. However, common ordering tasks such as sorting, ranking, or top-k selection are inherently non-differentiable due to their piecewise constant nature. This typically prevents the direct application of gradient descent, which is essential for efficient learning from data. These challenges have become increasingly relevant in recent years (Lapin et al., 2016; Blondel et al., 2020; Petersen et al., 2022b). To address them, relaxations or approximations are needed to make such tasks compatible with neural network training. This can be achieved by smoothing the objective functions, introducing uncertainty into algorithms, softening constraints, or adding variability. The known methods include smoothed approximations (Berrada et al., 2018; Garcin et al., 2022), optimal based (Cuturi et al., 2019; Xie et al., 2020), permutations (Petersen et al., 2022b). Some of these solutions are often not closed form.
@@
Challenges in addressing order related tasks arise from several factors. Among them are abrupt changes and instabilities when traversing regions. Time and memory issues often limit the solutions for large $n$ and $k$ values, where the aim lies in computing soft differenitiable top-k for large sequences of size $n$. There might also issues with scalability, parallel and batch processing. Techniques employed to solve these problems include relaxations and estimators, ranking regularizers, even learning-based ranking solutions. However, some problems remain unsolved or the existing methods are incomplete. E.g., efficiency in terms of time and memory usage may still be inadequate, impeding the solution of certain tasks (Xie et al., 2020). Frequently, there is a speed-precision trade-off. Several methods lack closed-form solutions, confining one to approximations or complicated calculations. Some methods do not yield probabilities (Blondel et al., 2020). Several do not yet support easy-to-use GPU methods.
Our aim was to address the above-mentioned issues. The main achievement of this paper is the construction of a general theory, based on arbitrary density, such that incorporation of the Laplace distribution provides a closed-form, numerically stable, and efficient solutions to all previously mentioned problems. More precisely, our approach, called $\mathrm{LapSum}$, is based on a sum of Laplace distributions, see the scheme Fig. 1 for construction of soft top-k selection problem. We show that this approach gives a solution that is both theoretically sound and practical: the reduced $O(n\log{}n)$ time complexity is on par with that of sequence sorting, while the memory requirements are very limited in comparison to other available top-k solutions.
@@
  • propose a novel, theoretically sound simple closed-form solution, called $\mathrm{LapSum}$, for all classical soft-order based problems [^1] Our basic top-k solution’s pseudo-code is just 26 lines long (see Appendix A).,
  • prove and experimentally verify that $\mathrm{LapSum}$ has $O(n\log{}n)$ low time and $O(n)$ memory complexities, together with derivatives, outperforming within this aspect all other existing methods, see  Fig. 2 for comparison with other SOTA methods,
  • offer easy to use code for $\mathrm{LapSum}$, for both CPU and CUDA, which makes our approach feasible for large optimization problems thanks to efficient use of parallelization. [^2] Full code is available at github.com/gmum/LapSum.
Figure 2: Upper: the relationship between the data dimension $n$ (horizontal axis) and the maximum memory usage and computation time (vertical axis), with $k=n/2$ in this figure. Computations for forward/backward processes shown here were performed on a CPU (see Appendix Figs. 11, 13, 10 and 12 for more). Below: the critical confidence diagram showing the statistical confidence of the above results (Demšar, 2006). $\mathrm{LapSum}$ approach is comparable to the best few (better on the right, statistically comparable joined with horizontal segments).
2 Related Work
The strategy of direct top-k training to optimize network performance, rather than straightforward use of top-1, was first introduced in fast top-k SVM (Lapin et al., 2016). The proposition called for the construction of differentiable loss functions to enable gradient learning.
Several methodologies are available. Some focus on smooth approximations. Berrada et al. (2018) proposed a smooth top-k to more accurately approximate the top-k loss and facilitate a direct technique for top-k gradient learning. Although these and comparable methods for direct gradient-based top-k learning yield precise results, they can be computationally intensive (Garcin et al., 2022). Some smooth approximations can speed up optimization, but may sacrifice accuracy.
@@ 2 Related Work
Multiple studies demonstrate effective applications of top-k learning techniques in pattern recognition and feature extraction, the creation of sparse networks and feature extraction, and recommender systems, such as through the use of differentiable logic gates (Zhao et al., 2019; Hoefler et al., 2021; Chen et al., 2023; Xu et al., 2023; Chen et al., 2021).
@@ 3 General soft-order theory based on sums of cumulative density functions
We begin by demonstrating how a cumulative density function can be used to construct soft analogues of hard, order-based problems such as top-k selection and permutations. Although the proposed approach is primarily theoretical and not computationally efficient for a general density, in the next section we show that it can be efficiently computed with complexity $O(n\log n)$ in the specific case of the Laplace distribution.
@@ 3.1 Function $F\text{-}\mathrm{Sum}$}
We base our construction on an arbitrary even and strictly positive density function $f$, with $F$ denoting its cumulative density function, so that F(x)=1F(x)F(-x)=1-F(x) and consequently $F(0)=1/2$. Additionally, for a scale parameter $\alpha\neq 0$ we put
@@
${F\text{-}\mathrm{Sum}}_{\alpha}(r,x)=\sum_{i}F_{\alpha}(x-r_{i}).$
Clearly, the image of ${F\text{-}\mathrm{Sum}}_{\alpha}$ is the interval $(0,n)$. It occurs that ${F\text{-}\mathrm{Sum}}_{\alpha}(r,x)$ is invertible as a function of $x$. This follows from the fact that since $f$ is a strictly positive density, $F_{\alpha}$ is a strictly increasing continuous function for $\alpha>0$ and strictly decreasing for $\alpha<0$. Consequently, ${F\text{-}\mathrm{Sum}}_{\alpha}(x,r)$ is a continuous strictly monotonous function, which is concluded in the following observation.
@@
Let sequence $r=(r_{0},\ldots,r_{n-1})\in\mathbb{R}^{n}$ be given, and let $k\in(0,n),\alpha\neq 0$ be arbitrary. Then the equation
${F\text{-}\mathrm{Sum}}_{\alpha}(r,x)=k$
has a unique solution $x$, which we denote by $x={F\text{-}\mathrm{Sum}}^{-1}(k)$
It should be noted, that in our reasoning we do not require $k$ to be integer.
@@
Note that the above functions are differentiable and that the derivatives can be computed efficiently. By $f_{\alpha}$ we denote the function $f_{\alpha}(x)=\frac{1}{\alpha}f(x/\alpha)$. Observe that, for $\alpha>0$, it is the density whose CDF is equal to $F_{\alpha}$. Then we have
$\frac{\partial}{\partial x}{F\text{-}\mathrm{Sum}}_{\alpha}(r,x)=\sum_{i}f_{\alpha}(x-r_{i}).$ (1)
Now we compute the derivatives of the inverse function $k\to{F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(r,k)$. Given $k\in(0,n)$, let $b=b_{\alpha}(r,k)$ denote the unique solution to
${F\text{-}\mathrm{Sum}}(b)=\sum_{i}{}F_{\alpha}(b-r_{i})=k.$
To compute the derivative of the function $b$ with respect to $w$, we differentiate the above formula with respect to $k$ obtaining $\frac{\partial b}{\partial k}\sum_{i}f_{\alpha}(b-r_{i})=1$. Thus
$\frac{\partial}{\partial k}{F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(r,k)\;\;=\frac{\partial b}{\partial k}\;\;=\frac{1}{\sum_{i}f_{\alpha}(b-r_{i})}.$ (2)
@@ 3.2 Soft rankings and soft orderings
Figure 3: The computation of the soft ranking based on $F$ given as CDF of Laplace distribution, here $(s_{0},s_{1},s_{2})={F\text{-}\mathrm{Rank}}_{\alpha=1}(r_{0},r_{1},r_{2})$.
For $\alpha\neq 0$, fixed sequence $(r_{i})$ and an arbitrary $j\in\{0,\ldots,n-1\}$ we define
${F\text{-}\mathrm{Rank}}_{\alpha}(r_{j})={F\text{-}\mathrm{Sum}}_{\alpha}(r,r_{j})-\tfrac{1}{2},$
see Fig. 3. By a soft ranking of whole sequence $r=(r_{i})$ we understand
${F\text{-}\mathrm{Rank}}_{\alpha}(r)=({F\text{-}\mathrm{Rank}}_{\alpha}(r_{i}))_{i}.$
First, we show that ${F\text{-}\mathrm{Rank}}_{\alpha}(r_{j})\subset(0,n-1).$ This follows directly from the formula
${F\text{-}\mathrm{Rank}}_{\alpha}(r_{j})=\sum_{l\neq j}F_{\alpha}(r_{j}-r_{l}),$ (3)
which is a consequence of the equality $F_{\alpha}(r_{j}-r_{j})=F(0)=\tfrac{1}{2}$.
One can easily observe that ${F\text{-}\mathrm{Rank}}_{\alpha}$ is permutation-equivariant, that is, a permutation of the input will result in a similar permutation of the output. By $s^{+}_{i}$ we denote the rank of $r_{i}$ in sequence $r$ (with respect to standard order), i.e.
@@
With $s_{i}^{-}$ we denote the rank of element $r_{i}$ in the reverse order of $r$, that is, $s_{i}^{-}=(n-1)-s_{i}^{+}$. To prove that ${F\text{-}\mathrm{Rank}}_{\alpha}$ is a correct soft version of ranking operation, we need to show that in the limiting case $\alpha\to 0^{\pm}$ we obtain the hard ranking.
@@
${F\text{-}\mathrm{Rank}}_{\alpha}(r_{j})\to s_{j}^{\pm}\text{ as }\alpha\to 0^{\pm},$
for arbitrary $j\in\{0,\ldots,n-1\}$.
@@
Let us prove the limiting formula above for $\alpha>0$ (the case $\alpha<0$ is similar). Since $F_{\alpha}$, for $\alpha\to 0^{+}$, converges to a Heaviside function, by (3) we get
$\lim\limits_{\alpha\to 0^{+}}{F\text{-}\mathrm{Rank}}_{\alpha}(r_{j})$ $=\lim\limits_{\alpha\to 0^{+}}\sum\limits_{i\neq j}F_{\alpha}(r_{j}-r_{i})$
$=\sum\limits_{i\neq j}H(r_{j}-r_{i})=s_{j}^{+}.\hskip 0.09995pt\qed$
Thus, with $\alpha\to 0^{+}$, we obtain the hard ranking of the sequence $r$ in increasing order, while for $\alpha\to 0^{-}$ we obtain the ranking in decreasing order.
@@
Soft ordering can be intuitively seen as an inverse operator to soft ranking. Given a sequence $(r_{i})$ we define
$({F\text{-}\mathrm{Sort}}_{\alpha}(r))_{l}={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(\tfrac{1}{2}+l)\text{ for }l\in\{0,\ldots,n-1\}.$
Informally we can write
${F\text{-}\mathrm{Sort}}_{\alpha}(r)={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}\left(\tfrac{1}{2}+(0,1,\ldots,n-1)\right).$
@@
  • it is permutation invariant,
  • ${F\text{-}\mathrm{Sort}}_{\alpha}(r)$ converges to the sorted (with respect to increasing order) $r$ for $\alpha\to 0^{+}$,
  • ${F\text{-}\mathrm{Sort}}_{\alpha}(r)$ converges to sorted $r$ (with respect to decreasing order) as $\alpha\to 0^{-}$.
Figure 4: The result of soft sorting with $F$ given by CDF of Laplace distribution. Observe that for small $\alpha$ the constructed sequence $(q_{0},q_{1},q_{2})={F\text{-}\mathrm{Sort}}_{\alpha=1}(r_{0},r_{1},r_{2})$ practically coincides with the sorted sequence $r$.
@@ 3.3 Soft top-k
We proceed to the construction of soft differentiable version of top-k selection problem. The aim is to construct a differentiable version of the hard top-k (formally max top-k) operation which, for a sequence $r=(r_{i})\in\mathbb{R}^{n}$, returns a binary sequence $(p_{i})\in\{0,1\}^{n}$ that has a value $1$ for the indices with k highest values of $r$, and zero for all other indices. Clearly, $\sum_{i}p_{i}=k$.
In our approach, we can choose an arbitrary $k\in(0,n)$, where we underline that we do not require $k$ to be an integer. Using $F\text{-}\mathrm{Sum}$} we construct the parameterization of the space
$\Delta_{k}=\{[p_{0},\ldots,p_{n-1}]\in(0,1)^{n}:\sum_{i}p_{i}=k\}.$
@@
$\begin{array}[]{c}{F\text{-}\mathrm{Top}}_{\alpha}(r,k)=(F_{\alpha}(b-r_{0}),\ldots,F_{\alpha}(b-r_{n-1}))\\ \text{ where }b={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k),\end{array}$
see Fig. 1 for a visualization of this approach.
@@
For every $k\in(0,n)$ and $r\in\mathbb{R}^{n}$ we have
${F\text{-}\mathrm{Top}}_{\alpha}(r,k)\in\Delta_{k}.$ (4)
Moreover, if $k$ is integer and the elements of $r$ are pairwise distinct, then
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r,k)\to\mathop{\mathrm{topmin}}_{k}(r)$ as $\alpha\to 0^{+}$,
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r,k)\to\mathop{\mathrm{topmax}}_{k}(r)$ as $\alpha\to 0^{-}$.
@@
Observe that (4) follows directly from the fact that $b={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k)$, and thus $\sum_{i}F_{\alpha}(b-r_{i})=k$.
Due to the limited space, the proof of the limiting case is given in the Appendix Theorem B.1. ∎
@@
One can observe that the function $F\text{-}\mathrm{Top}$} has properties similar to that of softmax. In particular, if two outputs are equal, then the inputs are equal modulo some constant translation. Namely, by the invertibility of $F$ we obtain that if ${F\text{-}\mathrm{Top}}_{\alpha}(r,k)={F\text{-}\mathrm{Top}}_{\alpha}(\bar{r},k)$ then
@@
where $b={F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(r,k)$, $\bar{b}={F\text{-}\mathrm{Sum}}_{\alpha}^{-1}(\bar{r},k)$. This implies, in particular, that similarly to softmax, ${F\text{-}\mathrm{Top}}_{\alpha}$ with domain restricted to $\{r=(r_{i})\in\mathbb{R}^{n}:\sum_{i}r_{i}=0\}$ is invertible.